# A theory of phonon-induced friction on molecular adsorbates

### v2

@@ 4.1.2 Surface desorption
The formalism described in the previous subsection can be applied to derive a rate constant for surface desorption. For gas-phase desorption, the reaction coordinate can be defined as the distance of the adsorbate center of mass from the surface. Furthermore, in many cases, the desorption process is barrierless, meaning that at the transition state the molecule and surface do not interact 50. Thus, we can adapt Eq. 25 to express the gas-phase desorption rate constant,
@@
While transition state theory is primarily a classical theory 51, 52, 53, 54, studies have demonstrated improved agreement with experiment when introducing quantum corrections, such as accounting for the rotational motion of the molecule or using quantum harmonic oscillator partition functions instead of classical oscillator partition functions 50, 55. We will thus compare four different models for the desorption rate constant to experimental results: (1) a fixed-surface model using classical harmonic oscillator partition functions and a rotation correction,
@@ 4.2 Results and discussion
We have compared results from Eqs. 30 to  33 to experimental temperature-dependent desorption rate constants for CO and Xe from a Pt(111) surface. Parameters used in computing Eqs. 30 to  33 are presented in Table 1. Note that because Xe is an atomic species, its moment of inertia is 0, and thus we ignore the rotational partition function factors for Xe calculations. The phonon corrections in Eq. 31 and Eq. 33 were computed using a 4x4x8 EMT surface slab and subsequently averaged across the first Brillouin zone to approximate an infinite surface, as described in Sec. 3. In the SI, we illustrate that the rate constant corrections we present here are not sensitive to the size of the surface slab used, or whether one accounts for phonon dispersion.
Table 1: Parameters used for computing desorption rate constants shown in Figure 5.
$E^{{\ddagger}}$ ($\mathrm{eV}$)
$\omega_{\mathrm{as}}$ ($\mathrm{cm^{-1}}$)
$\tilde{\omega}_{\mathrm{as}}$ ($\mathrm{cm^{-1}}$)
$\tilde{\omega}_{\mathrm{D}}$ ($\mathrm{cm^{-1}}$)
CO
1.4755
48032
164
203
Xe
0.24556
2834
21
156
The CO desorption rate constants were taken from Ref. 55 and the Xe desorption rates were taken from Ref. 56. In Ref. 55, desorption rate constants were calculated by fitting the time-dependent flux from a beam scattering experiment to two models: a single exponential model and a bi-exponential model. The single exponential model fit the entire flux signal, mixing contributions from terrace and steps. Meanwhile the biexponential model separated the flux into a fast component, arising from terrace desorption, and a slow component, arising from step to terrace diffusion followed by terrace desorption. While our TST calculations do not include the role of steps, we compare the results of our models to data both from the single exponential model and the fast component of the biexponential for thoroughness and transparency.
@@
In Fig. 5B, we demonstrate results for Xe. Here, all the TST models lie essentially on top of each other and are lower in value than the experimental rate constants, although by a small margin. The smaller phonon corrections for Xe versus CO are a natural result of the weaker interactions with the surface. Weaker coupling means a small $K(t=0)$, leading to $\tilde{\omega}_{\mathrm{as}}\approx\omega_{\mathrm{as}}$. Furthermore, weak coupling also results in a phonon density of states that is unchanged from a bare lattice, implying $\tilde{\omega}_{\mathrm{D}}=\omega_{\mathrm{D}}$. In general, the theory presented in this section suggests the phonon corrections to the rate constant are much larger for chemisorbed species than physisorbed species.
@@ 5 Methods
All calculations were performed using the Atomic Simulation Environment (ASE) python package 57, 58, 59, 60, 61, 62. Mass-weighted Hessians were computed using a finite-difference scheme in the vibrations module of ASE. Dynamical matrices were computed using the phonon module. The displacement size used in finite difference calculations was 0.01. 100 points were used to uniformly sample the first Brillouin zone.

### v3

@@ 4.1.2 Surface desorption
The formalism described in the previous subsection can be applied to derive a rate constant for surface desorption. For gas-phase desorption, the reaction coordinate can be defined as the distance of the adsorbate center of mass from the surface. Furthermore, in many cases, the desorption process is barrierless, meaning that at the transition state the molecule and surface do not interact 53. Thus, we can adapt Eq. 25 to express the gas-phase desorption rate constant,
@@
While transition state theory is primarily a classical theory 54, 55, 56, 57, studies have demonstrated improved agreement with experiment when introducing simple quantum corrections, such as accounting for the rotational motion of the molecule or using quantum harmonic oscillator partition functions instead of classical oscillator partition functions 53, 58. We will thus compare four different models for the desorption rate constant to experimental results: (1) a fixed-surface model using classical harmonic oscillator partition functions and a rotation correction,
@@ 4.2 Results and discussion
We have compared results from Eqs. 30 to  33 to experimental temperature-dependent desorption rate constants for CO and Xe from a Pt(111) surface. Parameters used in computing Eqs. 30 to  33 are presented in Table 1. We ignore the rotational partition function factors for Xe calculations. The phonon corrections in Eq. 31 and Eq. 33 were computed using a 4x4x8 EMT surface slab and subsequently averaged across the first Brillouin zone to approximate an infinite surface, as described in Sec. 3. In the SI, we illustrate that the rate constant corrections we present here are not sensitive to the size of the surface slab used, or whether one accounts for phonon dispersion.
Table 1: Parameters used for computing desorption rate constants shown in Fig. 5.
$E^{{\ddagger}}$ ($\mathrm{eV}$) $\omega_{\mathrm{as}}$ ($\mathrm{cm^{-1}}$) $\tilde{\omega}_{\mathrm{as}}$ ($\mathrm{cm^{-1}}$) $\tilde{\omega}_{\mathrm{D}}$ ($\mathrm{cm^{-1}}$)
CO 1.4758 48034 164 203
Xe 0.24559 2836 21 156
The CO desorption rate constants were taken from Ref. 58 and the Xe desorption rates were taken from Ref. 59. In Ref. 58, desorption rate constants were calculated by fitting the time-dependent flux from a beam scattering experiment to two models: a single exponential model and a bi-exponential model. The single exponential model fit the entire flux signal, mixing contributions from terrace and steps. Meanwhile the biexponential model separated the flux into a fast component, arising from terrace desorption, and a slow component, arising from step to terrace diffusion followed by terrace desorption. While our TST calculations do not include the role of steps, we compare the results of our models to data both from the single exponential model and the fast component of the biexponential for thoroughness and transparency.
@@
In Fig. 5B, we demonstrate results for Xe. Here, all the TST models lie essentially on top of each other and are lower in value than the experimental rate constants, although by a small margin. The smaller phonon corrections for Xe versus CO are a natural result of the weaker interactions with the surface. Weaker coupling means a small $K(t=0)$, leading to $\tilde{\omega}_{\mathrm{as}}\approx\omega_{\mathrm{as}}$. Furthermore, weak coupling also results in a phonon density of states that is unchanged from a bare lattice, implying $\tilde{\omega}_{\mathrm{D}}=\omega_{\mathrm{D}}$. In general, the theory presented in this section suggests the phonon corrections to the rate constant are much larger for chemisorbed species than physisorbed species. We also note that the deviation between theory and experiment in Fig. 5B is larger for lower temperatures. For sufficiently low-mass species at sufficiently cryogenic temperatures, it has been shown that tunneling plays a significant role in barrier crossing60, requiring more sophisticated quantum corrections to the rate constants than those in Eq. 32 and Eq. 33.
@@ 5 Methods
All calculations were performed using the Atomic Simulation Environment (ASE) python package 61, 62, 63, 64, 65, 66. Mass-weighted Hessians were computed using a finite-difference scheme in the vibrations module of ASE. Dynamical matrices were computed using the phonon module. The displacement size used in finite difference calculations was 0.01. 100 points were used to uniformly sample the first Brillouin zone.